# Scaling Towards the Information Boundary of Instruction Sets: The Infinity Instruct Subject Technical Report

### v3

@@ 6.2 Model Self-Improvement
Model self-improvement aims to iteratively enhance model capabilities through self-generated data or feedback signals. Representative approaches include self-refinement (madaan2023self; zhou2024cycle), multi-round generation and evaluation loops, and performance-driven data augmentation (yang2025knowing). These methods enable models to identify and address their own weaknesses by generating more challenging training samples, thereby facilitating continual improvement.
To ensure data quality, recent studies introduce deficiency diagnosis mechanisms (lighterness2024data), which analyze model performance on downstream tasks to detect knowledge gaps or skill deficiencies. These insights guide the targeted synthesis of training data(miller2025comparison). By incorporating such feedback into the training loop, models and datasets can co-evolve—improving both model performance and data coverage over time(bauer2024feedback).
Beyond simply increasing instruction quantity, recent research emphasizes instruction coverage (task/domain diversity) and depth (reasoning complexity). Several works have explored combining hierarchical tagging with complexity control to synthesize increasingly challenging and diverse instructions via evolutionary algorithms (lu2023instag; zhao2025beyond). These strategies enhance the cognitive depth of training data and promote models’ abilities in complex reasoning and compositional generalization.

### v4

@@ 6.2 Model Self-Improvement
Model self-improvement aims to iteratively enhance model capabilities through self-generated data or feedback signals. Representative approaches include self-refinement (Madaan et al., 2023; Zhou et al., 2024), multi-round generation and evaluation loops, and performance-driven data augmentation (Yang et al., 2025). These methods enable models to identify and address their own weaknesses by generating more challenging training samples, thereby facilitating continual improvement.
To ensure data quality, recent studies introduce deficiency diagnosis mechanisms (Lighterness et al., 2024), which analyze model performance on downstream tasks to detect knowledge gaps or skill deficiencies. These insights guide the targeted synthesis of training data(Miller et al., 2025). By incorporating such feedback into the training loop, models and datasets can co-evolve—improving both model performance and data coverage over time(Bauer et al., 2024).
Beyond simply increasing instruction quantity, recent research emphasizes instruction coverage (task/domain diversity) and depth (reasoning complexity). Several works have explored combining hierarchical tagging with complexity control to synthesize increasingly challenging and diverse instructions via evolutionary algorithms (Lu et al., ; Zhao et al., 2025). These strategies enhance the cognitive depth of training data and promote models’ abilities in complex reasoning and compositional generalization.